A plane flies 5000 km in 6.5 hours. How far would the plane fly in 10 hours?
Answer by jim_thompson5910(35256)

(5000 km)/(6.5 hours) = (x km)/(10 hours)


5000/6.5 = x/10


5000*10 = 6.5*x


50000 = 6.5x


50000/6.5 = x


7,692.30769 = x


x = 7,692.30769


So it traveled approximately 7,692.30769 km